# Holistic Memory Diversification for Incremental Learning in Growing Graphs

### v1

@@ Abstract
This paper addresses the challenge of incremental learning in growing graphs with increasingly complex tasks. The goal is to continually train a graph model to handle new tasks while retaining its inference ability on previous tasks. Existing methods usually neglect the importance of memory diversity, limiting in effectively selecting high-quality memory from previous tasks and remembering broad previous knowledge within the scarce memory on graphs. To address that, we introduce a novel holistic Diversified Memory Selection and Generation (DMSG) framework for incremental learning in graphs, which first introduces a buffer selection strategy that considers both intra-class and inter-class diversities, employing an efficient greedy algorithm for sampling representative training nodes from graphs into memory buffers after learning each new task. Then, to adequately rememorize the knowledge preserved in the memory buffer when learning new tasks, we propose a diversified memory generation replay method. This method first utilizes a variational layer to generate the distribution of buffer node embeddings and sample synthesized ones for replaying. Furthermore, an adversarial variational embedding learning method and a reconstruction-based decoder are proposed to maintain the integrity and consolidate the generalization of the synthesized node embeddings, respectively. Finally, we evaluate our model on node classification tasks involving increasing class numbers. Extensive experimental results on publicly accessible datasets demonstrate the superiority of DMSG over state-of-the-art methods.
@@ 1 Introduction
Graphs, owing to their flexible relational data structures, are widely employed for many applications in various domains, including social networks [1, 2], recommendation systems [3, 4], and bioinformatics [5]. With the increasing prevalence of graph data, graph-based models like Graph Neural Networks (GNNs) have gained significant attention due to their ability to capture complex structural relationships and those dynamic variants also demonstrate remarkable inductive capabilities on growing graph data [6, 7, 8]. However, as the growing graph in Figure 1 shows, when new nodes are added, the associated learning tasks can become increasingly complex. For example, the graph models on academic networks might need to predict the topics of papers in highly dynamic research areas where the topic rapidly emerges, and those on recommendation networks might need to continually adapt to new user preferences. Recent research on the inductive capability and adaptability of GNNs often remains limited to a specific task [9, 10, 11] and cannot be readily applied to incremental tasks. Moreover, it is often inefficient to train an entirely new model from scratch every time a new learning task is introduced. In recent years, model reuse via incremental learning [12, 13], also known as continual learning or lifelong learning, has led to exploring more economically viable pipelines, enabling the model to adaptively learn new tasks while maintaining the knowledge from old tasks.
@@
The main challenge of incremental learning on graphs lies in mitigating catastrophic forgetting. As the graph model learns from a sequence of tasks on evolving graphs, it tends to forget the information learned from previous tasks when acquiring knowledge from new tasks. One prevalent approach to address this issue is the memory replay method, a human-like method that typically maintains a memory buffer to store the knowledge gained from previous tasks. When learning a new task, the model not only focuses on the current information but also retrieves and re-learns from memory, preventing the model from forgetting what was learned previously as it takes on new tasks. This method has two major focuses to address on graphs: (1) How to select knowledge from old graphs to form more high-quality memory buffers? Existing methods usually select representative training samples as knowledge. However, determining which nodes in the graph are more representative is difficult and usually a time-consuming process. Furthermore, most methods [14, 15] select samples all into one same buffer without considering the inter-class differences between various previous tasks, which may degrade the quality of preserved knowledge. (2) How to effectively replay the limited buffer knowledge to enhance the model’s memorization of previous tasks? Due to constraints related to memory and training expenses, the samples chosen for the buffer are often limited. Many methods [16, 17] concentrate on memory selection, neglecting to broaden the boundaries of memory within the buffer, resulting in a discount in replay. Finding an effective way to replay knowledge from these limited nodes is critical to incremental learning in graphs.
In this paper, we propose a novel Diversified Memory Selection and Generation (DMSG) method on incremental learning in growing graphs, devised to tackle the above challenges. we consider that selecting diversified memory helps in Comprehensive Knowledge Retention: we apply a heuristic diversified memory selection strategy that takes into account both intra-class and inter-class diversities between nodes. By employing an efficient greedy algorithm, we selectively sample representative training nodes from the growing graph, placing them into memory buffers after completing each new learning task. Furthermore, we explore the memory diversification in memory reply for Enhanced Knowledge Memorization: we introduce a generative memory replay method, which first leverages a variational layer to produce the distribution of buffer node embeddings, from which synthesized samples are drawn for replaying. We incorporate an adversarial variational embedding learning technique and a reconstruction-based decoder. These are designed to preserve the integrity of the information and strengthen the generalization of the synthesized node embeddings on the label space, ensuring the essential knowledge is carried over accurately and effectively.
@@ 2 Problem Formulation
In this section, we present the formulation for the incremental learning problem in growing graphs. Generally, a growing graph is represented by a sequential of $m$ snapshots: $G=\{G^{1},G^{2},...,G^{m}\}$, and each snapshot corresponds to the inception of a new task, represented as $\mathcal{T}=\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{m}\}$. Each graph $G^{i}$ is evolved from the previous graph $G^{i-1}$, i.e., $G^{i-1}\subset G^{i},\forall i\in 2,...,m.$, and each learning task $\mathcal{T}^{i}$ is more complex than the previous task $\mathcal{T}_{i-1}$. This paper specifies the learning tasks to classification tasks, i.e., the number of classes increases alongside graph growth, increasing the task complexity. In this scenario, we aim to continually learn a model $f(\theta)$ on $\mathcal{T}$. For the $t$-th step, the task $\mathcal{T}^{t}$ incorporates a training node set $\mathcal{V}^{t}$ with previously unseen labels (i.e., novel classes), where each vertex $v_{i}\in\mathcal{V}^{t}$ has the label $y_{i}\in\mathcal{Y}^{t}$, and $\mathcal{Y}^{t}=\{y^{t}_{1},y^{t}_{2},...,y^{t}_{n}\}$ is the set of $n$ novel classes. The task $\mathcal{T}^{t}$ is to train the $f(\theta):v_{i}\rightarrow\bigcup_{j=1}^{t}\mathcal{Y}^{j}$ to ensure it can infer well on the current novel classes while preventing catastrophic forgetting of the inference ability on previous classes.
Jointly Incremental Learning. This is a straightforward solution for the problem, which collects training nodes of all classes of previous tasks to train $f(\theta)$ in each step. This treats the accumulated tasks $\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{t}\}$ as a whole new task and retrains the model from scratch. However, this solution is inefficient because it leads to redundant training of labeled nodes and creates computational challenges due to the growing graph size. Conversely, the buffer memory replay model offers a more practical solution.
Memory Replay for Incremental Learning. This method, instead of gathering all previous training nodes, maintains buffers $\mathcal{B}$ that store a small yet representative subset of training nodes for each class of previous tasks. The objectives can be formulated as follows:
@@
Figure 2: The framework of DMSG. In this instance, the graph model underwent training on a 2-class node classification task on $G^{1}$. Two new classes of nodes are added to form $G^{2}$. Certain nodes of the previous two classes are first selected into buffers. Then, the model is further trained on the two new classes of nodes and buffers to perform incremental learning.
@@ 3 Methodology
Without loss of generality, we choose the plain GCN model followed by a classifier head as the backbone of $f(\theta)$, which can encode each node $v_{i}$ into an embedding $z_{i}$ and a probability $p_{i}$. As shown in Figure 2, initially, this model is trained on the graph $G_{1}$. When the $t$-th task introducing new classes arrives, we first extend new parameters (highlighted as the yellow segment) into the last layer of the classifier, ensuring the output probabilities encompass the previous and newly introduced classes. To facilitate continual training of $f(\theta)$, we leverage the memory replay framework, which incorporates both the Heuristic Diversified Memory Selection (Section 3.1) and the Diversified Memory Generation Replay (Section 3.2).
@@
Since the $\Sigma$ measure the distribution diversity and $\mathcal{B}_{<t}$ is the subset of $\mathcal{G}_{<t}$ and is typically less diverse. Assuming the sampling strategy is unbiased upon means. As the $\mathcal{B}_{<t}$ more diversified, $\Sigma_{q}\rightarrow\Sigma_{p}$, leading to the Wasserstein distance decreases. Based on Theorem 1, the discrepancy between the expected loss under true distribution and the buffer distribution becomes less, making the optimization on the buffer more closely approximate the optimization on all previous graph data.
@@ 3.1 Heuristic Diversified Memory Selection
Heuristic Greedy Solution. However, achieving this objective for selecting different classes of buffers is an NP-hard problem. This kind of problem is usually addressed using heuristic methods [18]. Thus, we introduce a greedy algorithm to sample representative training nodes when new tasks are introduced. Specifically, suppose $\mathcal{V}^{t}=\bigcup_{i=K+1}^{K+n}\mathcal{C}^{t}_{i}$ is the training nodes of $t$-th task and $\mathcal{C}^{t}_{i}$ is the training set corresponding to the $i$-th novel class. We have previously selected buffers $\{\mathcal{B}_{i}\}_{i=1}^{K}$ in previous $t-1$ tasks, where $K$ and $n$ are the numbers of previous classes and novel classes, respectively. Then, the greedy selection strategy is defined in the Algorithm 1, where $\mathbb{D}(\mathcal{B}_{i})=\sum_{v\in\mathcal{B}_{i}}(\mathcal{A}(v,\mathcal{B}_{i})+\frac{1}{K-1}\sum_{j=1,j\neq i,\mathcal{B}_{j}\subset\mathcal{C}_{j}}^{K}\mathcal{A}(v,\mathcal{B}_{j}))$ is the set score function defined on the buffer set $\mathcal{B}_{i}$ of the $i$-th class, $\triangle_{\mathbb{D}}(v|\mathcal{B}_{i})$ is the gain of $f$ choosing $v$ into $\mathcal{B}_{i}$, and $v*$ is the chosen node using the greedy strategy. In greedy Algorithm 1, the core idea is to make the currently best choice of buffer nodes at every step, hoping to obtain the global optimal solution for the objective Eq.4 through this local optimal choice.
@@
The above Proposition can be derived from the Greedy Approximation Guarantee for Monotonic and Submodular Functions [19] (proof can be found in Theorem 1 in the Appendix), given that our function $f$ is both monotonic (from Lemma 1) and submodular (from Lemma 2). The greedy algorithm is guaranteed to produce a solution that is at least $(1-\frac{1}{e})$ times the optimal solution. ∎
@@ 3.2 Diversified Memory Generative Replay
Maintaining Integrity of Synthesized Embeddings. Then, we further propose to maintain the integrity of these variational node embeddings to prevent them from deviating too far, which implies that while the generated embeddings should exhibit diversity, they must remain similar to the ground-truth ones. Specifically, we adopt an adversarial learning strategy [20]. We introduce an auxiliary discriminator, $\mathcal{D}:\mathbb{R}^{h}\rightarrow\mathbb{R}^{1}$ with parameters $\phi_{d}$, tasked with distinguishing the original embeddings and the variational embeddings of nodes. In contrast, the model is learned to generate diversified variational node embeddings from the original ones, meanwhile ensuring the authenticity of synthesized variational node embeddings. Thus, the learning objective is defined as follows:
@@ 3.3 Overall Optimization
Synchronized Min-Max Optimizating. A gradient reversal layer (GRL) [21] is introduced between the variational embedding and the auxiliary discriminator so as to conveniently perform min-max optimization on $\theta,\phi_{v}$ and $\phi_{d}$ under $\mathcal{L}_{adv}$ in the same training step. GRL acts as an identity transformation during the forward propagation and changes the signs of the gradient from the subsequent networks during the backpropagation.
Scaliability on Large-Scale graphs. In practice, in each training step, training the proposed method on the entire graph and buffer at once may not be practical, especially for large-scale graphs. Following methods like GraphSAGE [9] and GraphSAINT [22], we adopt a mini-batch optimization strategy. We sample a multi-hop neighborhood for each node and set two kinds of batch sizes, $B^{\text{new}}$ and $B^{\text{buffer}}$, for the new task and replay losses in Eq. 11, respectively. The ratio of these batch sizes corresponds to the ratio between the total training nodes in the new task and the buffer.
@@ 4 Experiments
Experiment Setup. In this section, we describe the experiments we perform to validate our proposed method. We use the four growing graph datasets, CoraFull, OGB-Arxiv, Reddit, and OGB-Products, introduced in Continual Graph Learning Benchmark (CGLB) [23]. These graphs contain 35, 20, 20, and 23 sub-graphs, respectively, where each sub-graph corresponds to new tasks with novel classes. For baselines, we establish the upper bound baseline Joint defined in Section 2. The lower bound baseline Fine-tune employs only the newly arrived training nodes for model adaptation withour memory replay. Then, we set multiple continual learning models for graph as baselines, including EWC [24], MAS [25], GEM [26], LwF [27], TWP [28], ER-GNN [16], and SEM [14]. For a fair comparison, the backbone is set as two layers and the hidden dimension as 256 for all baselines. For evaluation metric, we first report the accuracy matrix $M^{acc}\in\mathbb{R}^{T\times T}$, which is lower triangular where $M^{acc}_{i,j}(i\geq j)$ represents the accuracy on the $j$-th tasks after learning the task $i$. To derive a single numeric value after learning all tasks, we report the Average Accuracy (AA) $\frac{1}{T}\sum_{i=1}^{T}M^{acc}_{T,i}$ and the Average Forgetting (AF) $\frac{1}{T-1}\sum_{i=1}^{T-1}M^{acc}_{T,i}-M^{acc}_{i,i}$ for each task after learning the last task. For detailed introduction to dataset description, compared methods, evaluation metrics, and hyperparameter setting, please refer to Appendix.
@@
Overall Comparison. This experiment aims to answer: How is DMSG’s performance on the continual learning on graphs? We compare DMSG with various baselines in the class-incremental continual learning task and report the experimental results in Table 1. Initially, we observe that DMSG attains a significant margin over other baseline methods across all datasets. Certain baseline methods demonstrate exceedingly poor results. This can be attributed to the difficulty of the problem, which involves more than 20 timesteps’ continual learning. When the model forgets intermediate tasks, errors are cumulatively compounded for subsequent tasks, potentially leading to the model easily collapsing. However, our model addresses this challenge through superior buffer selection and replay training strategies, effectively avoiding catastrophic problems. Among the various baselines, the most comparable method to DMSG is SEM. Our method outperforms SEM mainly because we improve the buffer selection strategy, i.e., instead of random selection, we employ distance measures to choose more representative nodes for each class. Additionally, the variational replay method enables our model to effectively learn the data distribution from previous tasks. When compared to Joint, DMSG achieves comparable results, demonstrating its effectiveness in preserving knowledge from previous tasks, even with limited training samples. Notably, our method outperforms Joint on the Reddit dataset, and also exhibits a positive AF. This improvement can be attributed to the proposed buffer selection strategy, which can select representative nodes, in other words, eliminate the noise nodes, and thereby enhancing the results.
@@
In-Depth Analysis of Continuous Performance.
This experiment aims to answer: How does DMSG’s fine-grained performance evolve after continuously learning each task? To present a more fine-grained demonstration of the model’s performance in continual learning on graphs, we analyzed the average performance across all previous tasks each time a new task was learned. The comparative results of Fine-tune, Joint, DMSG, and the top-performing baseline, SEM, are depicted in Figure 3. The curve represents the model’s performance after $t$ in terms of AA on all previous $t$ tasks. Also, we visualize the accuracy matrices of DMSG and SEM on the OGB-Arxiv and OGB-Product datasets. The results are presented in Figure 4. In these matrices, each row represents the performance across all tasks upon learning a new one, while each column captures the evolving performance of a specific task as all tasks are learned sequentially. In the visual representation, lighter shades signify better performance, while darker hues indicate inferior outcomes. From the results, we observed that as the number of tasks increases, the learning objectives grow increasingly complex, resulting in a reduction in performance across all examined methods, including Joint. That is because as tasks accumulate and the learning objectives become multifaceted, it becomes challenging for models to maintain optimal performance across all classes. Notably, the Fine-Tuning strategy experienced a substantial decline, with the model collapsing with the arrival of merely two new tasks, demonstrating that catastrophic forgetting occurs almost immediately when the model fails to access previous memories. This reinforces the need for effective continual learning techniques on the growing graphs where new tasks frequently emerge. While the performance drop was observed across all methods, DMSG demonstrated resilience and outperformed the top-performing baseline SEM. Also, DMSG predominantly displays lighter shades across the majority of blocks compared to SEM in Figure 4. Moreover, its competitive performance with Joint in specific datasets signifies its robustness and capability. This could be attributed to diversified memory selection and generation in DMSG that not only help in mitigating forgetting but also in adapting efficiently to new tasks.
@@
Component Analysis of Memory Replay. This experiment aims to answer: Are all the proposed memory replay technologies of DMSG have the claimed contribution to continual learning? To investigate the distinct contributions of the diversified memory generation replay method, we conducted an ablation study on it. We design three variant methods for DMSG to verify the effectiveness of adversarial synthesized embedding learning and the graph reconstruction optimization objective. w/o $\mathcal{L}_{MISE}$: This variant excludes the adversarial learning loss for maintaining the integrity of synthesized embeddings. ; w/o $\mathcal{L}_{CGSE}$: This variant excludes the graph reconstruction loss on variational embeddings for consolidating their generalization to label space; w/o all: both losses were removed, and as a result, the model operates without any variational embeddings and just replays the original nodes in the memory buffers. From the results in Table 2, we can observe when both $\mathcal{L}_{MISE}$ and $\mathcal{L}_{CGSE}$ are removed (w/o all), the performance is the lowest across all datasets. Also, a progressive improvement is observed as individual components in the model.

### v2

# Towards Continuous Reuse of Graph Models via Holistic Memory Diversification

@@ Abstract
This paper addresses the challenge of incremental learning in growing graphs with increasingly complex tasks. The goal is to continuously train a graph model to handle new tasks while retaining proficiency in previous tasks via memory replay. Existing methods usually overlook the importance of memory diversity, limiting in selecting high-quality memory from previous tasks and remembering broad previous knowledge within the scarce memory on graphs. To address that, we introduce a novel holistic Diversified Memory Selection and Generation (DMSG) framework for incremental learning in graphs, which first introduces a buffer selection strategy that considers both intra-class and inter-class diversities, employing an efficient greedy algorithm for sampling representative training nodes from graphs into memory buffers after learning each new task. Then, to adequately rememorize the knowledge preserved in the memory buffer when learning new tasks, a diversified memory generation replay method is introduced. This method utilizes a variational layer to generate the distribution of buffer node embeddings and sample synthesized ones for replaying. Furthermore, an adversarial variational embedding learning method and a reconstruction-based decoder are proposed to maintain the integrity and consolidate the generalization of the synthesized node embeddings, respectively. Extensive experimental results on publicly accessible datasets demonstrate the superiority of DMSG over state-of-the-art methods. The code and data can be found in https://github.com/joe817/DMSG.
@@ 1 Introduction
Graphs, owing to their flexible relational data structures, are widely employed for many applications in various domains, including social networks (Jiang et al., 2016; Qiao et al., 2021), recommendation systems (Wu et al., 2022), and bioinformatics (Zhang et al., 2021; Huang et al., 2025). With the increasing prevalence of graph data, graph-based models like Graph Neural Networks (GNNs) have gained significant attention due to their ability to capture complex structural relationships and those dynamic variants also demonstrate remarkable inductive capabilities on growing graph data (Ju et al., 2024; Pareja et al., 2020; Manessi et al., 2020). However, as the growing graph in Figure 1 shows, when new nodes are added, the associated learning tasks can become increasingly complex. For example, the graph models on academic networks might need to predict the topics of papers in highly dynamic research areas where the topic rapidly emerges, and those on recommendation networks might need to adapt continually to new user preferences. Recent research on the inductive capability and adaptability of GNNs often remains limited to a specific task (Hamilton et al., 2017; Han et al., 2021; Qiao et al., 2023) and cannot be readily applied to incremental tasks. Moreover, it is often inefficient to train an entirely new model from scratch every time a new learning task is introduced. In recent years, model reuse via incremental learning (Tan et al., 2022; Kim et al., 2022), also known as continual learning or lifelong learning, has led to exploring more economically viable pipelines, enabling the model to adaptively learn new tasks while maintaining the knowledge from old tasks.
The main challenge of incremental learning on graphs lies in mitigating catastrophic forgetting. As the graph model learns from a sequence of tasks on evolving graphs, it tends to forget the information learned from previous tasks when acquiring knowledge from new tasks. One prevalent approach to address this issue is the memory replay method, a human-like method that typically maintains a memory buffer to store the knowledge gained from previous tasks. When learning a new task, the model not only focuses on the current information but also retrieves and re-learns from memory, preventing the model from forgetting what was learned previously as it takes on new tasks. Nevertheless, this method has two major focuses to address on graphs: (1) How to select knowledge from old graphs to form more high-quality memory buffers? Existing methods usually select representative training samples as knowledge. However, determining which nodes in the graph are more representative is difficult and usually a time-consuming process. Furthermore, most methods (Zhang et al., 2022c; Wang et al., 2022) select samples all into one same buffer without considering the inter-class differences between various previous tasks, which may degrade the quality of preserved knowledge. (2) How to effectively replay the limited buffer knowledge to enhance the model’s memorization of previous tasks? Due to constraints related to memory and training expenses, the samples chosen for the buffer are often limited. Many methods (Zhou & Cao, 2021; Su & Wu, 2023) concentrate on memory selection, neglecting to broaden the boundaries of memory within the buffer, resulting in a discount in replay. Finding an effective way to replay knowledge from these limited nodes is critical to incremental learning in graphs.
@@
In this paper, we propose a novel Diversified Memory Selection and Generation (DMSG) method on incremental learning in growing graphs, devised to tackle the above challenges. We consider that selecting diversified memory helps in Comprehensive Knowledge Retention: we apply a heuristic diversified memory selection strategy that takes into account both intra-class and inter-class diversities between nodes. By employing an efficient greedy algorithm, we selectively sample representative training nodes from the growing graph, placing them into memory buffers after completing each new learning task. Furthermore, we explore the memory diversification in memory reply for Enhanced Knowledge Memorization: we introduce a generative memory replay method, which first leverages a variational layer to produce the distribution of buffer node embeddings, from which synthesized samples are drawn for replaying. We incorporate an adversarial variational embedding learning technique and a reconstruction-based decoder. These are designed to preserve the integrity and strengthen the generalization of the synthesized node embeddings on the label space, ensuring the essential knowledge is carried over accurately and effectively.
@@ 2 Problem Formulation
In this section, we present the formulation for the incremental learning problem in growing graphs. Generally, a growing graph is represented by a sequential of $m$ snapshots: $G=\{G^{1},G^{2},...,G^{m}\}$, and each snapshot corresponds to the inception of a new task, represented as $\mathcal{T}=\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{m}\}$. Each graph $G^{i}$ is evolved from the previous graph $G^{i-1}$, i.e., $G^{i-1}\subset G^{i},\forall i\in 2,...,m.$, and each learning task $\mathcal{T}^{i}$ is more complex than the previous task $\mathcal{T}_{i-1}$. This paper specifies the learning tasks to classification tasks and follows the class-incremental learning setting, i.e., the number of classes increases alongside graph growth, increasing the task complexity. In this scenario, we aim to continuously learn a model $f(\theta)$ on $\mathcal{T}$. For the $t$-th step, the task $\mathcal{T}^{t}$ incorporates a training node set $\mathcal{V}^{t}$ with previously unseen labels (i.e., novel classes), where each vertex $v_{i}\in\mathcal{V}^{t}$ has the label $y_{i}\in\mathcal{Y}^{t}$, and $\mathcal{Y}^{t}=\{y^{t}_{1},y^{t}_{2},...,y^{t}_{n}\}$ is the set of $n$ novel classes. The task $\mathcal{T}^{t}$ is to train the $f(\theta):v_{i}\rightarrow\bigcup_{j=1}^{t}\mathcal{Y}^{j}$ to ensure it can infer well on the current novel classes while preventing catastrophic forgetting of the inference ability on previous classes.
Jointly Incremental Learning. This is a straightforward solution for the problem, which collects training nodes of all classes of previous tasks to train $f(\theta)$ in each step. This treats the accumulated tasks $\{\mathcal{T}^{1},\mathcal{T}^{2},...,\mathcal{T}^{t}\}$ as a whole new task and retrains the model from scratch. However, this solution is inefficient because it leads to redundant training of labeled nodes and creates computational challenges due to the growing graph size.
Memory Replay for Incremental Learning. This method offers a more practical solution. Instead of gathering all previous training nodes, it maintains buffers $\mathcal{B}$ that store a small yet representative subset of training nodes for each class of previous tasks. The objectives can be formulated as follows:
@@
Figure 2: The framework of DMSG. In this instance, the graph model underwent training on a 2-class node classification task on $G^{1}$. Two new classes of nodes are added to $G^{1}$ to form $G^{2}$. Certain nodes of the previous two classes are first selected into buffers. Then, the model is further trained on the two new classes of nodes and buffers to perform incremental learning.
@@ 3 Methodology
Without loss of generality, we choose the plain GCN model followed by a classifier head as the backbone of $f(\theta)$, which encodes each node $v_{i}$ in graphs into an embedding $z_{i}$ and a probability $p_{i}$. As shown in Figure 2, initially, this model was trained on the graph $G_{1}$. When the $t$-th task introducing new classes arrives, we first extend new parameters (highlighted as the yellow segment) into the last layer of the classifier, ensuring the output probabilities encompass the previous and newly introduced classes. To facilitate continual training of $f(\theta)$, we leverage the memory replay framework, which incorporates both the Heuristic Diversified Memory Selection (Section 3.1) and the Diversified Memory Generation Replay (Section 3.2).
@@
Since the $\Sigma$ measure the distribution diversity and $\mathcal{B}_{<t}$ is the subset of $\mathcal{G}_{<t}$ and is typically less diverse (detailedly analyzed in Appendix A.1). Assuming the sampling strategy is unbiased upon means, as the $\mathcal{B}_{<t}$ more diversified, $\Sigma_{q}\rightarrow\Sigma_{p}$, leading to the Wasserstein distance decreases. Based on Theorem 1, the discrepancy between the expected loss under true distribution and the buffer distribution becomes less, making the optimization on the buffer more closely approximate the optimization on all previous graph data.
@@ 3.1 Heuristic Diversified Memory Selection
Heuristic Greedy Solution. However, achieving this objective for selecting different classes of buffers is an NP-hard problem. This kind of problem is usually addressed using heuristic methods (Hochbaum, 1996). Thus, we introduce a greedy algorithm to sample representative training nodes when new tasks are introduced. Specifically, suppose $\mathcal{V}^{t}=\bigcup_{i=K+1}^{K+n}\mathcal{C}^{t}_{i}$ is the training nodes of $t$-th task and $\mathcal{C}^{t}_{i}$ is the training set corresponding to the $i$-th novel class. We have previously selected buffers $\{\mathcal{B}_{i}\}_{i=1}^{K}$ in previous $t-1$ tasks, where $K$ and $n$ are the numbers of previous classes and novel classes, respectively. Then, the greedy selection strategy is defined in the Algorithm 1, where $\mathbb{D}(\mathcal{B}_{i})=\sum_{v\in\mathcal{B}_{i}}(\mathcal{A}(v,\mathcal{B}_{i})+\frac{1}{K-1}\sum_{j=1,j\neq i,\mathcal{B}_{j}\subset\mathcal{C}_{j}}^{K}\mathcal{A}(v,\mathcal{B}_{j}))$ is the set score function defined on the buffer set $\mathcal{B}_{i}$ of the $i$-th class, $\triangle_{\mathbb{D}}(v|\mathcal{B}_{i})$ is the gain of $f$ choosing $v$ into $\mathcal{B}_{i}$, and $v*$ is the chosen node using the greedy strategy. In greedy Algorithm 1, the core idea is to make the currently best choice of buffer nodes at every step, hoping to obtain the global optimal solution for the objective Eq.4 through this local optimal choice.
@@
The above Proposition can be derived from the Greedy Approximation Guarantee for Monotonic and Submodular Functions (Nemhauser et al., 1978) (proof can be found in Theorem 2 in the Appendix), given that our function $f$ is both monotonic (from Lemma 1) and submodular (from Lemma 2). The greedy algorithm is guaranteed to produce a solution that is at least $(1-\frac{1}{e})$ times the optimal solution. ∎
@@ 3.2 Diversified Memory Generative Replay
Maintaining Integrity of Synthesized Embeddings. Then, we further propose to maintain the integrity of these variational node embeddings to prevent them from deviating too far, which implies that while the generated embeddings should exhibit diversity, they must remain similar to the ground-truth ones. Specifically, we adopt an adversarial learning strategy. We introduce an auxiliary discriminator, $\mathcal{D}:\mathbb{R}^{h}\rightarrow\mathbb{R}^{1}$ with parameters $\phi_{d}$, tasked with distinguishing the original embeddings and the variational embeddings of nodes. In contrast, the model is learned to generate diversified variational node embeddings from the original ones, meanwhile ensuring the authenticity of synthesized variational node embeddings. Thus, the learning objective is defined as follows:
@@ 3.3 Overall Optimization
Synchronized Min-Max Optimization. A gradient reversal layer (GRL) (Ganin et al., 2016) is introduced between the variational embedding and the auxiliary discriminator so as to conveniently perform min-max optimization on $\theta,\phi_{v}$ and $\phi_{d}$ under $\mathcal{L}_{MISE}$ in the same training step. GRL acts as an identity transformation during the forward propagation and changes the signs of the gradient from the subsequent networks during the backpropagation.
Scaliability on Large-Scale graphs. In each training step, training DMSG on the entire graph and buffer at once may not be practical, especially for large-scale graphs. Following methods like GraphSAGE (Hamilton et al., 2017) and GraphSAINT (Zeng et al., 2019), we adopt a mini-batch optimization strategy. We sample a multi-hop neighborhood for each node and set two kinds of batch sizes, $B^{\text{new}}$ and $B^{\text{buffer}}$, for the new task and replay losses in Eq. 11, respectively. The ratio of these batch sizes corresponds to the ratio between the total training nodes in the new task and the buffer.
@@ 4 Experiments
Experiment Setup. In this section, we describe the experiments we perform to validate our proposed method. We use the four growing graph datasets, CoraFull, OGB-Arxiv, Reddit, and OGB-Products, introduced in Continual Graph Learning Benchmark (CGLB) (Zhang et al., 2022a). These graphs contain 35, 20, 20, and 23 sub-graphs, respectively, where each sub-graph corresponds to new tasks with novel classes. For baselines, we establish the upper bound baseline Joint defined in Section 2. The lower bound baseline Fine-tune employs only the newly arrived training nodes for model adaptation without memory replay. Then, we set multiple continual learning models for the graph as baselines, including EWC (Kirkpatrick et al., 2017), MAS (Aljundi et al., 2018), GEM (Lopez-Paz & Ranzato, 2017), LwF (Li & Hoiem, 2017), TWP (Liu et al., 2021), ER-GNN (Zhou & Cao, 2021), SSM (Zhang et al., 2022c), and SEM (Zhang et al., 2023). For a fair comparison, the backbone is set as two layers and the hidden dimension as 256 for all baselines. For evaluation metric, we first report the accuracy matrix $M^{acc}\in\mathbb{R}^{T\times T}$, which is lower triangular where $M^{acc}_{i,j}(i\geq j)$ represents the accuracy on the $j$-th tasks after learning the task $i$. To derive a single numeric value after learning all tasks, we report the Average Accuracy (AA) $\frac{1}{T}\sum_{i=1}^{T}M^{acc}_{T,i}$ and the Average Forgetting (AF) $\frac{1}{T-1}\sum_{i=1}^{T-1}M^{acc}_{T,i}-M^{acc}_{i,i}$ for each task after learning the last task. For a more detailed introduction to the experimental setup, please refer to Appendix A.5.
@@
Overall Comparison. This experiment aims to answer: How is DMSG’s performance on the continual learning on graphs? We compare DMSG with various baselines in the class-incremental continual learning task and report the experimental results in Table 1. Initially, we observe that DMSG attains a significant margin over other baseline methods across all datasets. Certain baseline methods demonstrate exceedingly poor results. This can be attributed to the difficulty of the problem, which involves more than 20 timesteps’ continual learning. When the model forgets intermediate tasks, errors are cumulatively compounded for subsequent tasks, potentially leading to the model easily collapsing. However, our model addresses this challenge through superior buffer selection and replay training strategies, effectively avoiding catastrophic problems. Among the various baselines, the most comparable method to DMSG is SEM. Our method outperforms SEM mainly because we improve the buffer selection strategy, i.e., instead of random selection, we employ distance measures to choose more representative nodes for each class. Additionally, the variational replay method enables our model to effectively learn the data distribution from previous tasks. When compared to Joint, DMSG achieves comparable results, demonstrating its effectiveness in preserving knowledge from previous tasks, even with limited training samples. Notably, our method outperforms Joint on the Reddit dataset, and also exhibits a positive AF. This improvement can be attributed to the proposed buffer selection strategy, which can select representative nodes, in other words, eliminate the noise nodes, and thereby enhance the results.
@@
In-Depth Analysis of Continuous Performance. This experiment aims to answer: How does DMSG’s fine-grained performance evolve after continuously learning each task? To present a more fine-grained demonstration of the model’s performance in continual learning on graphs, we analyzed the average performance across all previous tasks each time a new task was learned. The comparative results of Fine-tune, Joint, DMSG, and the top-performing baseline, SEM, are depicted in Figure 3. The curve represents the model’s performance after $t$ in terms of AA on all previous $t$ tasks. Also, we visualize the accuracy matrices of DMSG and SEM on the OGB-Arxiv and OGB-Product datasets. The results are presented in Figure 4. In these matrices, each row represents the performance across all tasks upon learning a new one, while each column captures the evolving performance of a specific task as all tasks are learned sequentially. In the visual representation, lighter shades signify better performance, while darker hues indicate inferior outcomes. From the results, we observed that as the number of tasks increases, the learning objectives grow increasingly complex, resulting in a reduction in performance across all examined methods, including Joint. That is because as tasks accumulate and the learning objectives become multifaceted, it becomes challenging for models to maintain optimal performance across all classes. Notably, the Fine-Tuning strategy experienced a substantial decline, with the model collapsing with the arrival of merely two new tasks, demonstrating that catastrophic forgetting occurs almost immediately when the model fails to access previous memories. This reinforces the need for effective continual learning techniques on the growing graphs where new tasks frequently emerge. While the performance drop was observed across all methods, DMSG demonstrated resilience and outperformed the top-performing baseline SEM. Also, DMSG predominantly displays lighter shades across the majority of blocks compared to SEM in Figure 4. Moreover, its competitive performance with Joint in specific datasets signifies its robustness and capability. This could be attributed to diversified memory selection and generation in DMSG that not only help in mitigating forgetting but also in adapting efficiently to new tasks.
@@
Component Analysis of Memory Replay. This experiment aims to answer: Are all the proposed memory replay technologies of DMSG have the claimed contribution to continual learning? To investigate the distinct contributions of the diversified memory generation replay method, we conducted an ablation study on it. We design three variant methods for DMSG— w/o $\mathcal{L}_{MISE}$: This variant excludes the adversarial learning loss for maintaining the integrity of synthesized embeddings; w/o $\mathcal{L}_{CGSE}$: This variant excludes the graph reconstruction loss on variational embeddings for consolidating their generalization to label space; w/o all: both losses were removed, and as a result, the model operates without any variational embeddings and just replays the original nodes in the memory buffers. From the results in Table 2, we can observe when both $\mathcal{L}_{MISE}$ and $\mathcal{L}_{CGSE}$ are removed (w/o all), the performance is the lowest across all datasets. Also, a progressive improvement is observed as individual components in the model. This confirms their respective contributions to continual learning. An interesting trend emerges when comparing the individual contributions of the two components.